OpenStudy (anonymous):

Can someone please explain how to solve this problem? I don't remember planes... In the xy-plane, what is the slope of the line whose equation is 3x-2y=8?

jimthompson5910 (jim_thompson5910):

First solve for y. Do you know how to do this?

OpenStudy (anonymous):

I think so, like this?: subtract 3x from each side: -2y=8-3x divide each side by -2 y=8-3x  -2

OpenStudy (anonymous):

um, that didn't come out right. I was trying to show that 8-3x is over -2

jimthompson5910 (jim_thompson5910):

like this \[\Large 3x-2y=8\] \[\Large -2y=8-3x\] \[\Large y=\frac{8-3x}{-2}\] ??

OpenStudy (anonymous):

precisely

jimthompson5910 (jim_thompson5910):

Keep going and simplify it into y = mx+b form

OpenStudy (anonymous):

THAT I don't know

jimthompson5910 (jim_thompson5910):

\[\Large 3x-2y=8\] \[\Large -2y=8-3x\] \[\Large y=\frac{8-3x}{-2}\] \[\Large y=\frac{-3x+8}{-2}\] \[\Large y=\frac{-3x}{-2}+\frac{8}{-2}\] \[\Large y=\frac{3}{2}x-4\] Notice how the last equation is in y = mx+b form where m = 3/2 and b = -4

OpenStudy (anonymous):

Ok, I guess that could make sense. But then what does that mean? I still don't understand where y is if I were to chart it on the graph

OpenStudy (anonymous):

and what does "slope of the line" mean?

jimthompson5910 (jim_thompson5910):

y = mx+b is the general equation of any line

jimthompson5910 (jim_thompson5910):

The value of m is the slope

jimthompson5910 (jim_thompson5910):

The point (0,b) is the y-intercept

OpenStudy (anonymous):

Ok, so 3/2 is the slope then?

jimthompson5910 (jim_thompson5910):

yes, this represents how "steep" or "shallow" the line is

jimthompson5910 (jim_thompson5910):

it also represents if the line goes up or down (as you read it from left to right)

OpenStudy (anonymous):

ok, so if I were to graph this, what would it look like?

jimthompson5910 (jim_thompson5910):

You would plot a point at (0,-4), which is the y-intercept

OpenStudy (phi):

Here is a video, on graphing http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/graphs-using-slope-intercept-form he also has other videos about lines that might be helpful

jimthompson5910 (jim_thompson5910):

Then you would go up 3 units, then to the right 2 units to get to the next point (2, -1)

jimthompson5910 (jim_thompson5910):

You now have two points. Draw a straight line through them to graph the line.

OpenStudy (anonymous):

Cool, thank you for all your help, jim! Also, thank you, phi for the vid link, I'll check it out :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (phi):

This one looks more appropriate for how to graph the line http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/algebra--graphing-lines-1